OpenStudy (anonymous):

Please help with all of them I will give a medal. 1. The data set below shows the weight (in pounds) of 5 different pumpkins. What is the mean of the following set of data? 7,5,9,5,12 A. 7.3 pounds B. 7.4 pounds C. 7.5 pounds D. 7.6 pounds

OpenStudy (anonymous):

Do you know how to find the mean?

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

Okay so go ahead, and figure out the mean of this set of numbers. When you finish that it's your answer. You need to figure something our yourself if you know how to do it.

OpenStudy (anonymous):

Well I just learned mean so I want to know if I was figuring it out right

OpenStudy (anonymous):

Oh okay! So what your going to do it add up all of the numbers. Once you finish that you will divide it by the amount of numbers there are. So in your case it would be 7+5+9+5+12=x (X= represents your answer) Now when you find the answer for that you will divide it by the amount of numbers there are In your case it would be x÷5= Your answer (There are 5 numbers in that set) Got it? :o)